whats the simpliest way to show the square root of 96
Answer by stanbon(75887)

whats the simpliest way to show the square root of 96
---
sqrt(96) = sqrt(16*6) = sqrt(15)*sqrt(6) = 4*sqrt(6)
=======================
cheers,
Stan H.
====================